Solve the system of equations
y^2-3x^2=6
y=2x-1

Answer by lwsshak3(3700)

Solve the system of equations
y^2-3x^2=6
y=2x-1
**
substitute (2x-1) for y in first equation
y^2-3x^2=6
(2x-1)^2-3x^2=6
4x^2-4x+1-3x^2=6
x^2-4x-5=0
(x-5)(x+1)=0
x=5
or
x=-1